# Retrieval Enhanced Feedback via In-context Neural Error-book

### v4

@@ Abstract
Jongyeop Hyun School of CSE Chung-Ang University jesussuej@cau.ac.krthanks: contact: mldljyh@postech.ac.kr          Bumsoo Kim School of CSE Chung-Ang University bumsoo@cau.ac.kr
@@ 1 Introduction
Recent LLM advancements show superior reasoning performance, with extensive research on in-context learning (ICL) to enhance human-like reasoning capabilities. Early works on ICL generated responses based on a few provided correct examples, enabling the models to adapt to new tasks without extensive retraining (Brown et al., 2020; Dong et al., 2024; Wei et al., 2023). Although ICL has demonstrated effectiveness by primarily leveraging correct examples (Min et al., 2022), subsequent works have re-examined the fact that human learning is also deeply rooted in learning from errors (Edmondson, 1996; Chialvo and Bak, 1999; Edmondson, 1999). Recent studies suggest that incorporating errors into the learning process can further improve LLM performance (Sun et al., 2024). These approaches typically identify recurring errors, extract underlying principles from them, and apply these insights to prevent similar errors in the future.
However, a critical limitation of these methodologies is the absence of a systematic framework for structuring errors, making it difficult to analyze and mitigate failure cases effectively. This challenge becomes even more pronounced in Multimodal Large Language Models (MLLMs) that jointly process multimodal inputs (Zhao et al., 2023). Unlike unimodal LLMs where errors in textual reasoning can often be traced and corrected through established interpretability techniques, MLLMs introduce additional complexity due to the integration of visual and textual modalities. Without a structured approach to diagnosing and addressing previous errors, failures in one modality can propagate through the system, making it more difficult to ensure reliable and interpretable outcomes (Lau et al., 2025). While recent works such as multimodal-CoT have shown that incorporating textual CoT reasoning and in-context learning can improve performance, a fundamental gap persists: the lack of structured error analysis for multimodal reasoning makes it unclear whether existing MLLMs can fully leverage CoT reasoning for visual understanding (Alayrac et al., 2022; Zhao et al., 2023). Furthermore, the interpretability of MLLMs remains a significant challenge, as current approaches do not adequately explain how visual information contributes to reasoning, highlighting the need for a more rigorous framework for structuring and mitigating errors in multimodal AI systems.
@@ 2.1 Chain-of-Thought Reasoning
Chain-of-Thought (CoT) reasoning enhances Large Language Model (LLM) problem-solving by breaking down tasks into intermediate logical steps, akin to human cognition (Wei et al., 2022). This structured approach has improved performance in mathematical, commonsense, and multimodal tasks. However, CoT’s efficacy depends on the model’s inherent reasoning and prompt quality. A key challenge is error propagation from incorrect intermediate steps, necessitating refinements like retrieval or error-correction mechanisms for improved robustness (Cao et al., 2023).
@@ 2.2 Language Instruction Understanding in Multimodal Tasks
Language instructions are vital for AI-user interaction in multimodal tasks, traditionally requiring precision. While recent Multimodal Large Language Models (MLLMs) better handle complex instructions, ambiguity remains a challenge (Fu et al., 2023; Fei et al., 2022). Efforts like WAFFLECLIP (Roth et al., 2023) and FUDD (Esfandiarpoor and Bach, 2024) address polysemy in image classification, and REPHRASE (Prasad et al., 2024) uses iterative prompting for visual question answering. However, many existing methods demand extensive interactions or predefined rules, hindering generalization and scalability across diverse tasks and models.
@@ 2.3 Complex Reasoning with Multimodal Large Language Models
MLLMs increasingly integrate visual and textual information (Caffagni et al., 2024), yet comprehending complex language in visual contexts remains difficult (Zhao et al., 2023; Kil et al., 2024). Approaches to enhance MLLM reasoning are either training-based, aligning models with image-text data, sometimes using synthetic data (Davidson et al., 2025), or non-training-based, like CoT, which simulates step-by-step reasoning (Pramanick et al., 2024; Kil et al., 2024). Training-based methods can be data-intensive and costly (Davidson et al., 2025), while non-training methods often presume strong pre-existing reasoning capabilities (Zhang et al., 2024a), limiting general-purpose intelligence development.
@@ 2.4 Systematic Structuring of MLLMs
Systematically structuring MLLMs with Retrieval-Augmented Generation (RAG) (Lewis et al., 2020) is crucial for performance and scalability. Studies like Wang et al. (2024) and VisRAG (Yu et al., 2024) show RAG’s benefits in multimodal contexts, while RAVEN (Rao et al., 2024) highlights gains in tasks like image captioning. This underscores the need for integrating retrieval modules for adaptability. Our work also advocates for refined feedback flows, drawing on structured techniques to filter unhelpful signals, akin to our concise “Error-book” concept. Existing methods, however, often face data curation challenges or domain-specific limitations. Our approach aims for a distilled feedback structure, enhancing MLLM resilience and extensibility with minimal overhead.
@@ 2.5 Advancements in Error-Driven Learning
Recent error-driven learning methods include LEAP (Zhang et al., 2024b), generating static principles from errors, and TRAN (Tong et al., 2024), maintaining rules to avoid past mistakes. RICP (Sun et al., 2024) clusters errors into task-level principles and retrieves question-level insights. These methods often rely on generalized principles or pre-clustered errors, potentially misaligning with immediate task goals. Our approach differs by restructuring feedback based on the Feedback Model (Hattie and Timperley, 2007), focusing on task/process-level guidance through an “Error-book” addressing: Feed-Target (goal), Feed-Check (progress), and Feed-Path (actions). This prioritizes task-specific guidance, enabling efficient, deterministic retrieval without teacher intervention during inference, thus reducing overhead and improving precision over methods with rigid clustering or insufficient task focus.
@@ 3.1 Structured Feedback Generation
REFINE is a teacher-student framework where the teacher model systematically analyzes errors made by the student model. We designed three guiding principles to construct a final structured feedback for each error (student misprediction) inspired by classic educational psychology (Hattie and Timperley, 2007): Feed-Target, Feed-Check, and Feed-Path. These principles guide structured error analysis and facilitate the generation of precise, actionable insights. Given a multimodal QA benchmark, our process begins by evaluating the student model on a training set of image-question pairs $\{(I_{i},Q_{i})\}_{i=1}^{N}$. Errors from student predictions $\hat{A}_{i}$ (incorrect answers) are aggregated and provided to the teacher model alongside their ground-truth reference $A_{i}$. The teacher model generates structured feedback through the threefold analysis:
@@ 3.2 Feedback Filtering
After Feed-Target, Feed-Check and Feed-Path, the resulting feedback instances are categorized. This classification is performed automatically by our teacher model. The model is provided with the definitions of “Task/Process-relative” feedback (that directly corrects task-specific errors or adds specifics for reasoning, e.g., “Adjust counts for occluded objects.”) and “Self-Regulatory” feedback (that addresses metacognitive habits or personal traits, e.g., “To improve accuracy, try solving similar problems multiple times.”) from  Hattie and Timperley (2007) and then prompted to classify each generated instance. This automated method ensures the process is consistent and reproducible. Based on empirical observations suggesting that self-regulatory feedback tends to hinder final Chain-of-Thought (CoT) performance, we filter out feedback classified as self-regulatory.
@@ 3.3 Neural Error-book Construction
After filtering the generated feedback to retain only actionable task/process-level feedback (denoted as $\{\mathcal{F}_{i}\}_{i\in N_{e}}$, where $N_{e}$ denotes error cases), we pair each feedback with the joint embedding for the corresponding image-question pairs to construct a structured Neural Error-book database $R$. Given $\phi$ as a pre-trained embedding model (e.g., voyage-multimodal-3 (VoyageAI, 2024)) that obtains a joint multimodal embedding from each image-question pair in the training set $x_{i}=(I_{i},Q_{i})$, we construct a Neural Error-book $R$ as:
@@ 4.1 Experimental Setup
We evaluated the effectiveness of our feedback mechanisms using three multimodal reasoning benchmarks: MME-RealWorld (Zhang et al., 2024c), MMStar (Chen et al., 2024), and SEED-Bench-2-Plus (Li et al., 2024). These benchmarks collectively emphasize diverse multimodal reasoning capabilities, including error diagnosis, procedural correction, and text-rich visual comprehension. The evaluation framework employed for these benchmarks was VLMEvalKit (Duan et al., 2024). All models in our experiments were used with a temperature setting of 0.0, and results were reported using the pass@1 metric. Specifically:
@@
  • Pixtral-12B (Agrawal et al., 2024) (a high-capacity model)
  • Qwen2.5-VL-3B-Instruct (Bai et al., 2025) (a compact model)
Feedback generation employed Gemini-1.5-Pro (Team et al., 2024) for MME-RealWorld and MMStar, while SEED-Bench-2-Plus utilized Gemini-2.0-Flash (Pichai et al., 2024) due to its specialized text-rich processing capabilities.
@@
  • Standard Prompting (Brown et al., 2020): The LLM is asked to output the answer directly, without the intermediate reasoning process.
  • Chain of Thought (Wei et al., 2022): The LLM is instructed to think step-by-step before providing the answer.
  • Direct Feedback (Daheim et al., 2024): unstructured, open-ended feedback without explicit instructional framing.
  • RICP (Sun et al., 2024): retrieved principles from errors clustered by error type, providing task- and question-specific guidance.
@@ 4.3 Ablation Study: Analyzing Feedback Component Contributions
  • Cluster-level feedback: To assess generalized feedback via clustering (cf. RICP (Sun et al., 2024)), we applied K-means (k=5) to the multimodal embeddings of training instances. For each data cluster, our teacher model generated a single generalized feedback from 20 samples of task/process feedback within that cluster. At inference, a query was assigned to its nearest cluster (by embedding similarity), combining this generalized feedback with REFINE’s Task/Process feedback.
  • CoT (Wei et al., 2022): Standard Chain-of-Thought prompting (‘Let’s think step by step.’) was appended after REFINE’s Task/Process feedback to elicit explicit reasoning from the student model, rather than altering the feedback content itself.
  • Self-Reg (Hattie and Timperley, 2007): Self-regulatory feedback, normally filtered out (Section 3.2), was re-introduced alongside Task/Process feedback.

### v5

@@ Abstract
Jongyeop Hyun School of CSE Chung-Ang University jesussuej@cau.ac.krthanks: contact: mldljyh@postech.ac.kr          Bumsoo Kimthanks: Corresponding author. School of CSE Chung-Ang University bumsoo@cau.ac.kr
@@ 1 Introduction
Recent LLM advancements show superior reasoning performance, with extensive research on in-context learning (ICL) to enhance human-like reasoning capabilities. Early works on ICL generated responses based on a few provided correct examples, enabling the models to adapt to new tasks without extensive retraining (brown.etal_2020; dong-etal-2024-survey; wei.etal_2023a). Although ICL has demonstrated effectiveness by primarily leveraging correct examples (min-etal-2022-rethinking), subsequent works have re-examined the fact that human learning is also deeply rooted in learning from errors (Edmondson1996LearningFM; chialvo.bak_1999; Edmondson1999PsychologicalSA). Recent studies suggest that incorporating errors into the learning process can further improve LLM performance (sun-etal-2024-retrieved). These approaches typically identify recurring errors, extract underlying principles from them, and apply these insights to prevent similar errors in the future.
However, a critical limitation of these methodologies is the absence of a systematic framework for structuring errors, making it difficult to analyze and mitigate failure cases effectively. This challenge becomes even more pronounced in Multimodal Large Language Models (MLLMs) that jointly process multimodal inputs (zhao.etal_2023a). Unlike unimodal LLMs where errors in textual reasoning can often be traced and corrected through established interpretability techniques, MLLMs introduce additional complexity due to the integration of visual and textual modalities. Without a structured approach to diagnosing and addressing previous errors, failures in one modality can propagate through the system, making it more difficult to ensure reliable and interpretable outcomes (lau2025uncertainty). While recent works such as multimodal-CoT have shown that incorporating textual CoT reasoning and in-context learning can improve performance, a fundamental gap persists: the lack of structured error analysis for multimodal reasoning makes it unclear whether existing MLLMs can fully leverage CoT reasoning for visual understanding (alayrac.etal_2022a; zhao.etal_2023a). Furthermore, the interpretability of MLLMs remains a significant challenge, as current approaches do not adequately explain how visual information contributes to reasoning, highlighting the need for a more rigorous framework for structuring and mitigating errors in multimodal AI systems.
@@ 2.1 Chain-of-Thought Reasoning
Chain-of-Thought (CoT) reasoning enhances Large Language Model (LLM) problem-solving by breaking down tasks into intermediate logical steps, akin to human cognition (wei2022chain). This structured approach has improved performance in mathematical, commonsense, and multimodal tasks. However, CoT’s efficacy depends on the model’s inherent reasoning and prompt quality. A key challenge is error propagation from incorrect intermediate steps, necessitating refinements like retrieval or error-correction mechanisms for improved robustness (cao.etal_2023).
@@ 2.2 Language Instruction Understanding in Multimodal Tasks
Language instructions are vital for AI-user interaction in multimodal tasks, traditionally requiring precision. While recent Multimodal Large Language Models (MLLMs) better handle complex instructions, ambiguity remains a challenge (fu.etal_2023; fei-etal-2022-cqg). Efforts like WAFFLECLIP (roth2023waffling) and FUDD (esfandiarpoor2024follow) address polysemy in image classification, and REPHRASE (prasad2024rephrase) uses iterative prompting for visual question answering. However, many existing methods demand extensive interactions or predefined rules, hindering generalization and scalability across diverse tasks and models.
@@ 2.3 Complex Reasoning with Multimodal Large Language Models
MLLMs increasingly integrate visual and textual information (caffagni-etal-2024-revolution), yet comprehending complex language in visual contexts remains difficult (zhao.etal_2023a; NEURIPS2024_32923dff). Approaches to enhance MLLM reasoning are either training-based, aligning models with image-text data, sometimes using synthetic data (davidson.etal_2025), or non-training-based, like CoT, which simulates step-by-step reasoning (NEURIPS2024_d74033a2; NEURIPS2024_32923dff). Training-based methods can be data-intensive and costly (davidson.etal_2025), while non-training methods often presume strong pre-existing reasoning capabilities (zhang.etal_2024k), limiting general-purpose intelligence development.
@@ 2.4 Systematic Structuring of MLLMs
Systematically structuring MLLMs with Retrieval-Augmented Generation (RAG) (lewis2020retrieval) is crucial for performance and scalability. Studies like wang-etal-2024-searching and VisRAG (yu.etal_2024) show RAG’s benefits in multimodal contexts, while RAVEN (rao.etal_2024) highlights gains in tasks like image captioning. This underscores the need for integrating retrieval modules for adaptability. Our work also advocates for refined feedback flows, drawing on structured techniques to filter unhelpful signals, akin to our concise “Error-book” concept. Existing methods, however, often face data curation challenges or domain-specific limitations. Our approach aims for a distilled feedback structure, enhancing MLLM resilience and extensibility with minimal overhead.
@@ 2.5 Advancements in Error-Driven Learning
Recent error-driven learning methods include LEAP (zhang.etal_2024i), generating static principles from errors, and TRAN (tong.etal_2024a), maintaining rules to avoid past mistakes. RICP (sun-etal-2024-retrieved) clusters errors into task-level principles and retrieves question-level insights. These methods often rely on generalized principles or pre-clustered errors, potentially misaligning with immediate task goals. Our approach differs by restructuring feedback based on the Feedback Model (hattie.timperley_2007), focusing on task/process-level guidance through an “Error-book” addressing: Feed-Target (goal), Feed-Check (progress), and Feed-Path (actions). This prioritizes task-specific guidance, enabling efficient, deterministic retrieval without teacher intervention during inference, thus reducing overhead and improving precision over methods with rigid clustering or insufficient task focus.
@@ 3.1 Structured Feedback Generation
REFINE is a teacher-student framework where the teacher model systematically analyzes errors made by the student model. We designed three guiding principles to construct a final structured feedback for each error (student misprediction) inspired by classic educational psychology (hattie.timperley_2007): Feed-Target, Feed-Check, and Feed-Path. These principles guide structured error analysis and facilitate the generation of precise, actionable insights. Given a multimodal QA benchmark, our process begins by evaluating the student model on a training set of image-question pairs $\{(I_{i},Q_{i})\}_{i=1}^{N}$. Errors from student predictions $\hat{A}_{i}$ (incorrect answers) are aggregated and provided to the teacher model alongside their ground-truth reference $A_{i}$. The teacher model generates structured feedback through the threefold analysis:
@@ 3.2 Feedback Filtering
After Feed-Target, Feed-Check and Feed-Path, the resulting feedback instances are categorized. This classification is performed automatically by our teacher model. The model is provided with the definitions of “Task/Process-relative” feedback (that directly corrects task-specific errors or adds specifics for reasoning, e.g., “Adjust counts for occluded objects.”) and “Self-Regulatory” feedback (that addresses metacognitive habits or personal traits, e.g., “To improve accuracy, try solving similar problems multiple times.”) from  hattie.timperley_2007 and then prompted to classify each generated instance. This automated method ensures the process is consistent and reproducible. Based on empirical observations suggesting that self-regulatory feedback tends to hinder final Chain-of-Thought (CoT) performance, we filter out feedback classified as self-regulatory.
@@ 3.3 Neural Error-book Construction
After filtering the generated feedback to retain only actionable task/process-level feedback (denoted as $\{\mathcal{F}_{i}\}_{i\in N_{e}}$, where $N_{e}$ denotes error cases), we pair each feedback with the joint embedding for the corresponding image-question pairs to construct a structured Neural Error-book database $R$. Given $\phi$ as a pre-trained embedding model (e.g., voyage-multimodal-3 (voyage_2024d)) that obtains a joint multimodal embedding from each image-question pair in the training set $x_{i}=(I_{i},Q_{i})$, we construct a Neural Error-book $R$ as:
@@ 4.1 Experimental Setup
We evaluated the effectiveness of our feedback mechanisms using three multimodal reasoning benchmarks: MME-RealWorld (zhang.etal_2024j), MMStar (chen.etal_2024c), and SEED-Bench-2-Plus (li.etal_2024d). These benchmarks collectively emphasize diverse multimodal reasoning capabilities, including error diagnosis, procedural correction, and text-rich visual comprehension. The evaluation framework employed for these benchmarks was VLMEvalKit (duan2024vlmevalkit). All models in our experiments were used with a temperature setting of 0.0, and results were reported using the pass@1 metric. Specifically:
@@
  • Pixtral-12B (agrawal2024pixtral12b) (a high-capacity model)
  • Qwen2.5-VL-3B-Instruct (bai2025qwen25vltechnicalreport) (a compact model)
Feedback generation employed Gemini-1.5-Pro (geminiteam2024gemini15unlockingmultimodal) for MME-RealWorld and MMStar, while SEED-Bench-2-Plus utilized Gemini-2.0-Flash (gemini2.0flash_2024e) due to its specialized text-rich processing capabilities.
@@
  • Standard Prompting (brown.etal_2020): The LLM is asked to output the answer directly, without the intermediate reasoning process.
  • Chain of Thought (wei2022chain): The LLM is instructed to think step-by-step before providing the answer.
  • Direct Feedback (daheim-etal-2024-stepwise): unstructured, open-ended feedback without explicit instructional framing.
  • RICP (sun-etal-2024-retrieved): retrieved principles from errors clustered by error type, providing task- and question-specific guidance.
@@ 4.3 Ablation Study: Analyzing Feedback Component Contributions
  • Cluster-level feedback: To assess generalized feedback via clustering (cf. RICP (sun-etal-2024-retrieved)), we applied K-means (k=5) to the multimodal embeddings of training instances. For each data cluster, our teacher model generated a single generalized feedback from 20 samples of task/process feedback within that cluster. At inference, a query was assigned to its nearest cluster (by embedding similarity), combining this generalized feedback with REFINE’s Task/Process feedback.
  • CoT (wei2022chain): Standard Chain-of-Thought prompting (‘Let’s think step by step.’) was appended after REFINE’s Task/Process feedback to elicit explicit reasoning from the student model, rather than altering the feedback content itself.
  • Self-Reg (hattie.timperley_2007): Self-regulatory feedback, normally filtered out (Section 3.2), was re-introduced alongside Task/Process feedback.